in the problems x=4y+1 and x=4y-1 how do i solve for y. do not understand how to solve for y. thank you.
Answer by Alan3354(69443)

in the problems x=4y+1 and x=4y-1 how do i solve for y. do not understand how to solve for y.
--------------
x=4y+1 --> y = (x-1)/4
x=4y-1 --> y = (x+1)/4
------------------------
There is no solution to this system.
x cannot equal both 4y+1 and 4y-1
If you graph them, they're parallel lines. No intersection, no solution.